[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun Jul 30, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Jul 31 02:57:10 UTC 2017


HSU Shift Log for Sun Jul 30 2017 22:57:04 GMT-0400 (EDT)

Shift: R

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues: orange warning banner won't go away on the backup desk  
AWIPS2.

Other Techincal issues: HMON failed at 18z for AL98, according to the SDM.  
He also mentioned the upgrade date has slipped to August 15.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170731/75ef2850/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list